David Ortiz Motors has a target capital structure of 40% debt and 60% equity. The yield to maturity on the company's outstanding bonds is 12%, and the company's tax rate is 40%. Ortiz's CFO has calculated the company's WACC as 9.49%. What is the company's cost of equity capital? Round your answer to two decimal places.
Answer #1

WACC is weighted Avg cost of sources in capital structure.

Let X be the Cost of Equity.

After Tax cost of Debt = YTM (1- Tax Rate )

= 12% ( 1 - 0.4)

= 12% * 0.6

= 7.2%

WACC:

Source Weight Cost Wtd Cost
Debt 0.4 7.20% 0.0288
Equity 0.6 X 0.6X
WACC 0.0288 + 0.6X

Thus 0.0288 + 0.6X = 0.0949

0.6X = 0.0949 - 0.0288

= 0.0661

X = 0.661 / 0.6

= 0.1102 i.e 11.02%

Cost of Equity is 11.02%